# When in Doubt, Cascade: Towards Building Efficient and Capable Guardrails

### v1

@@ 1 Introduction
Large language models (LLMs) contain high potential for a variety of real-world applications, due to their versatility, adaptability, and ease of use, along with their continuously improving performance [29, 6, 28, 32]. Yet, their deployment, especially in critical domains such as healthcare and finance, poses significant risks [5, 18]. A new host of challenges arises with the generative capabilities of these models, as they may produce convincing output, but this output may often be layered with issues around toxicity, bias, hallucinations, and more.
In order to combat the harmful generations from these models, the concept of guardrail (or detector) models have gained popularity for several reasons. These guardrail models can be more efficient, modular, and scalable [1, 33, 19] than the LLMs whose output they operate on. In this work, we focus on the problem of detecting whether an LLM’s textual output contains social bias.
@@
Social bias can be defined as discrimination for, or against, a person or group, or a set of ideas or beliefs, in a way that is prejudicial or unfair [41, 7]. Importantly, text which contains social bias may not contain any explicit or profane content, but may still propagate discrimination (e.g., “I don’t want to hire this individual as a babysitter because they have facial scars."). Driven by a clear and present need to automatically detect whether LLM-generated text contains such harmful content, we developed a social-bias-detector. To do so, we gathered a collection of open source datasets, with commercially permissible licenses, and used a combination of four datasets as an attempt at a holistic collection of training data. We provide the specific datasets, as well as the hyper-parameters used during training, in Appendix A.1. From an architectural standpoint, the social-bias-detector is an encoder-only model with just over 100M parameters that was obtained by fine-tuning BERT [10].
Despite reasonable performance on our evaluation sets (specific numbers in Appendix A.2), we discovered a high false positive issue with this model. Here, a false positive refers to the classification of benign text as harmful. In order to investigate further, we devised an experiment to test the hypothesis that there existed a mismatch between the training dataset (which were largely human-generated or curated) and the distribution of text that is generated by an LLM. Three of the authors manually annotated outputs from the llama-2-7b-chat111https://huggingface.co/meta-llama/Llama-2-7b-chat-hf model and we conjectured that our model’s sub-optimal performance was due to the excessively intricate and evasive answers, which tend to be generated by highly aligned and verbose models such as Llama 2 [38]. However, concurrent to our work, we came across the true reason for our detector’s subpar performance - the use-mention distinction [14]. We will revisit this paper later in our experimental results, in Section 3.
In the context of social bias detection, the use-mention distinction can be thought of as the difference between using text for ill-intent and simply mentioning text without this malicious undertone. To elaborate, the text “All Muslims are terrorists" is an example of use, whereas the text “The notion that all Muslims are terrorists is a harmful stereotype" is an example of mention. For further contextualization of the use-mention distinction, as well as the impact of this phenomenon on downstream social bias detection, refer to Section 6. Extended analysis revealed that many of misclassifications from our detector were the result of failing to differentiate use from mention. In particular, a substantial quantity of the responses from the Llama-2 model were of the following flavor: “It is not accurate to say that {toxic_statement}," which precisely maps to a mention.
@@ 2.1 Taxonomy-Driven Synthetic Data Generation
We opted to utilize taxonomy driven generation, as this allowed us to explore a rich and structured manner through which to generate data at scale. First, we generated and iteratively updated a taxonomy of social biases. We started by prompting the Mixtral mixture-of-experts model222https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1, under greedy decoding, with the instruction: Generate a high level taxonomy for social bias that shows various groups and associated biases and stereotypes. After post-processing, we were able to arrive at a viable taxonomy with around 50 social biases across a variety of dimensions. Some examples include Anti-Asian bias, Transphobia, Ethnocentrism (the belief that one’s own culture is superior to others), etc. The full taxonomy that we utilized is available in Appendix B. Importantly, the taxonomy provides over 10 high-level topics (e.g., Racial and Ethnic Bias, Gender Bias, Classism, etc.), and a variety of sub-topics for each of these groups. In particular, we were encouraged to see that our generated taxonomy included demographic groups that are not as mainly focused on, but are still extremely relevant for the purposes hate speech detection [27]. One such example is Polyphobia, or prejudice against individuals who are attracted to multiple partners.
Once we generated this taxonomy, we proceeded with tailored synthetic data generation. Here, we directly leveraged the taxonomy and generated data for each leaf node (which is each sub-topic in our taxonomy). Using the prompt in Figure 1, we repeatedly instructed the aforementioned Mixtral mixture-of-experts model to generate pairs of output - jointly generating one biased and one unbiased statement. For reproducibility, we generated under nucleus sampling [17] with top_p at 0.95, top_k at 100, temp at 0.7, and max_new_tokens set to 1024. This method was desirable for a couple of reasons. First, the generation directly utilized the taxonomy, as each call for a generation contained both the root node (i.e., the over-arching topic) as well as the leaf node (i.e., the specific bias) directly in the context for the prompt. Next, requiring JSON formatting in the output allowed for easy post-processing of the LLM-generated text. Finally, and perhaps most interestingly, asking for paired output resulted in the generation of contrastive data. Specifically, we observed that each pair of {biased, unbiased} text precisely mapped to {use, mention} examples! Refer to Table 1 for the flavor of synthetic data obtained from this procedure.
@@ 2.2 Our Suite of Social-Bias-Detectors
social-bias refers to the models trained on the four human curated datasets mentioned in Section 2 and elaborated upon in Appendix A.1. In particular, we have three such models - each trained on top of a different base model. social-bias is trained on top of BERT333https://huggingface.co/google-bert/bert-base-uncased, social-bias-distil is trained on top of a transformer architecture [39] which provides most of the accuracy of a BERT-like model [10], while being seven times faster on a CPU and two times faster on a GPU. Finally, social-bias-toxigen is trained on top of toxigen_hatebert444https://huggingface.co/tomh/toxigen_hatebert[15]. Note that these three models were trained with only human curated datasets.
@@
We hypothesized that the cascade would be a “best-of-both-worlds" approach where would do a decent job on aggregate, due to the high quality human curated datasets that it has seen. Then, to precisely combat the use-mention distinction issue, where would incorrectly flag text that is a mention as harmful, we will run any harm-labeled instances from through , a model that is specifically trained, by way of our synthetic data, to distinguish use from mention.
@@ 3.1 Evaluation Datasets and Baselines
Due to the fact that our primary axis for evaluation is the use-mention distinction, we sought to construct an evaluation set which reflected this goal. Even though datasets for such a task are not widely available [14], we leveraged the rich literature on counter-speech and counter-narratives in the context of hate speech. In particular, similar to [14], we take the Knowledge-grounded hate countering [9] and the Multi-Target CONAN [13] datasets. The knowledge-grounded dataset contains 195 hate speech and counter-narrative pairs covering multiple hate targets (islamophobia, misogyny, antisemitism, racism, and homophobia). The counter narratives are written by an expert who is tasked with composing a suitable counter-narrative response to a given hate speech using the corresponding knowledge as much as possible [9]. The Multi-Target CONAN dataset consists of 5003 hate Speech and counter-narrative pairs covering multiple hate targets, including disabled, Jews, LGBT+, migrants, Muslims, people of color (POC), women. The dataset is constructed using a novel human-in-the-loop data collection methodology [13].
We combined both datasets and labeled each pair of {hate speech, counter narrative} as {bias, not bias}. In addition, we noted that a label of bias most precisely meant a label of use and similarly for not bias and mention. This is because counter-narratives are written such that they directly counteract the harmful hate speech [14], which results in them being excellent examples of mentions.
@@
We provide two sets of evaluations below. First, in order to compare with the results from [14], we utilize the same evaluation set of 180 total examples (also taken from the above two datasets), including both hate speech and counter-narratives. Second, we combine the entirety of the Knowledge-grounded hate countering and Multi-Target CONAN datasets to arrive at an evaluation set of size 555This comes from combining 195 pairs with 5003 pairs from the Knowledge-grounded hate countering and Multi-Target CONAN datasets, respectively..
@@
In addition to the reported numbers with three of the GPT models from [14], we also report numbers using Llama-Guard666https://huggingface.co/meta-llama/LlamaGuard-7b and Llama-Guard-2777https://huggingface.co/meta-llama/Meta-Llama-Guard-2-8B. We use these two models as both a competitive baseline as well as a point of comparison due to the size and latency of these models. To reiterate, the detectors are on the order of 100M parameters, Llama-Guard models are either 7B or 8B parameters, and GPT-4 is rumored to be orders of magnitude larger [30].
@@ 4 Experimental Results
In this section, we provide results on the aforementioned datasets, with all of our detectors and baselines. We report three metrics of interest: false positive rate (FPR), false negative rate (FNR), and average error rate (Avg Err) - similar to [14]. For clarification, the average error rate is the average of the FPR and FNR. To further contextualize these metrics, a false positive refers to incorrectly providing a label of bias to benign text. Moreover, because all of the examples of benign text are examples of mentions and all the examples of harmful text are uses, a false positive represents incorrectly flagging a mention as use. Symmetrically, a false negative refers to incorrectly flagging a use as mention. For our purposes, we note that both false positives and false negatives are important. False positives represent improper moderation by flagging benign text whereas false negatives represent missed detection by failing to flag harmful text.
@@ 5.1 Results for 180 Examples Test Set
First, we comment on results for the evaluation set of 180 examples from [14]. Despite the small size of this test, we wanted to demonstrate performance because [14] provided us with points of comparison to the GPT-family of models. In particular, we observe that our cascade models provide competitive performance with the GPT models. We make particular note of cascade-onetrial. Recall that the model which determines use or mention for cascade-onetrial has only seen one iteration of synthetic data generation, making it extremely desirable, as the burden of training data is substantially reduced. In absolute terms, for cascade-onetrial saw around 1K examples in training, whereas for cascade-distil saw around 190K examples in training (refer to Appendix C) for full details). Therefore, we are able to see that for this (albeit limited) set of data, the cascade approach is able to perform on par with some of the most widely used and largest models, while having a fraction of the computational cost. We do notice that both Llama-Guard models surpass even GPT-4 by a significant margin. Nevertheless, for this set of 180 examples, the cascade approach takes under 1 minute on a single A100 GPU, and is even able to be run on a CPU (taking a few hours). In comparison, access to GPT models is limited to querying via expensive external APIs, and inference with Llama-Guard models is not possible on a CPU.
@@ 5.3 Extended Analysis on Llama-Guard-2
Ignoring the increased computational overhead, we further explore the differences in performance (with respect to false positives) when comparing our detectors with the Llama-Guard models. At an instance level, we explored the differences in predictive quality between Llama-Guard and our detectors. For example, the sample Did you know that children with Down syndrome, if surrounded by regularly-developed peers, receive a fundamental boost in their development? This "mix" fosters a positive learning atmosphere and increases engagement and attention in all students. This mean that children with Down syndrome would be a burden to society if isolated, not the contrary. was incorrectly flagged as unsafe by Llama-Guard-2, but properly classified as not_bias by our cascade-orig method. Another such sample is Even if all British Muslims had cancer, you are safe: you cannot catch it from anyone! Cancer cannot be spread, since it is a non contagious disease. Also, it is worth considering that illnesses are not connected to religions.
@@ 6 Related Work
Social bias can be defined as discrimination for, or against, a person or group, or a set of ideas or beliefs, in a way that is prejudicial or unfair [41, 7]. Numerous studies have demonstrated that generative models exhibit undesirable behavior that amplifies social bias [3, 31, 36, 35, 11, 27]. Additionally, the deficiencies of current datasets [4] and opacity of defining what constitutes social bias in language models and their measures [3, 35, 1] have emerged.
@@
As discussed briefly at the end of Section 1, the use-mention distinction can be thought of as the difference between using text for ill-intent and simply mentioning text without this malicious undertone. Although a subtle difference, this phenomenon is essential for many downstream applications. For example, a significant amount of content online falls under the umbrella of mentions, such as counter-speech, media reporting, education, and legal settings [14, 25, 22, 16, 42]. Of particular use to us is the notion of counter-speech, which refers to speech produced by users of online platforms to counteract harmful speech of others [14]. By definition, counter-speech is an example of mention, and thus detection systems which are unable to distinguish use from mention risk contributing to downstream harm - such as improper removal of counter-speech. This in turn reduces opportunities to rectify false narratives and risks further censoring those already most affected by harmful language [14]. Recently, there has also been work which documents the real experiences of counter-speakers and the ways in which to address existing barriers [26]. An exciting by-product of our work was our own introduction to this rich space, and we believe that the area of counter-speech will continue to provide essential help in the plight of detecting and mitigating harmful content produced by generative models.
@@
There is a growing set of rich literature on guardrail models and how they can be more efficient, modular, and scalable [1, 33, 19]. NeMo Guardrails [33] is an open-source toolkit for adding modular and programmable guardrails to LLM-based conversational systems. Llama-Guard [19] is an LLM-based input-output safeguard model which has a customizable taxonomy of harms. Finally, the authors in [1] go over insights from developing a host of guardrail-like models, which they refer to as detectors.
@@
Central to our work is the leveraging of capable large language models to generate synthetic data which can be used to train our guardrail models. In particular, we leverage taxonomy-guided data generation. Previous work such as GLAN [23] utilizes a pre-curated taxonomy of human knowledge and capabilities as input and generates large-scale synthetic instruction data. However, as pointed out by [37], this method relies using the proprietary GPT-4 [30] as the teacher model, which imposes restrictions on the downstream commercial usability of the generated data. Concretely, the terms of use of proprietary models obfuscate the viability of training commercially viable models using data that is generated from proprietary or otherwise closed models - which many times forbid using their model to improve other models 999See point v under License Rights in https://llama.meta.com/llama3/license/. Thus, similar to [37], we utilize the open source Mixtral101010https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 model to generate our data. Additionally, the taxonomy guides the generation of the data, as it enables targeted coverage around the individual leaf nodes of the taxonomy [37].

### v2

@@ 1 Introduction
Large language models (LLMs) contain high potential for a variety of real-world applications, due to their versatility, adaptability, and ease of use, along with their continuously improving performance (OpenAI 2022; Bommasani 2023; Nayak 2019; Perspective API 2021). Yet, their deployment, especially in critical domains such as healthcare and finance, poses significant risks (IBM AI Risk Atlas ). A new host of challenges arises with the generative capabilities of these models, as they may produce convincing output, but this output may often be layered with issues around toxicity, bias, hallucinations, and more.
In order to combat the harmful generations from these models, the concept of guardrail (or detector) models have gained popularity for several reasons. These guardrail models can be more efficient, modular, and scalable (Achintalwar et al. 2024; Rebedea et al. 2023; Inan et al. 2023) than the LLMs whose output they operate on. In this work, we focus on the problem of detecting whether an LLM’s textual output contains social bias.
@@
Social bias can be defined as discrimination for, or against, a person or group, or a set of ideas or beliefs, in a way that is prejudicial or unfair (Webster et al. 2022; Bommasani and Liang 2022). Importantly, text which contains social bias may not contain any explicit or profane content, but may still propagate discrimination (e.g., “I don’t want to hire this individual as a babysitter because they have facial scars.”). Driven by a clear and present need to automatically detect whether LLM-generated text contains such harmful content, we developed a social-bias-detector. To do so, we gathered a collection of open source datasets, with commercially permissible licenses, and used a combination of four datasets as an attempt at a holistic collection of training data. We provide the specific datasets, as well as the hyper-parameters used during training, in Appendix A.1. From an architectural standpoint, the social-bias-detector is an encoder-only model with just over 100M parameters that was obtained by fine-tuning BERT (Devlin et al. 2019).
Despite reasonable performance on our evaluation sets (specific numbers in Appendix A.2), we discovered a high false positive issue with this model. Here, a false positive refers to the classification of benign text as harmful. In order to investigate further, we devised an experiment to test the hypothesis that there existed a mismatch between the training dataset (which were largely human-generated or curated) and the distribution of text that is generated by an LLM. Three of the authors manually annotated outputs from the llama-2-7b-chat111https://huggingface.co/meta-llama/Llama-2-7b-chat-hf model and we conjectured that our model’s sub-optimal performance was due to the excessively intricate and evasive answers, which tend to be generated by highly aligned and verbose models such as Llama 2 (Touvron et al. 2023). However, concurrent to our work, we came across the true reason for our detector’s subpar performance - the use-mention distinction (Gligoric et al. 2024). We will revisit this paper later in our experimental results, in Section 3.
In the context of social bias detection, the use-mention distinction can be thought of as the difference between using text for ill-intent and simply mentioning text without this malicious undertone. To elaborate, the text “All Muslims are terrorists” is an example of use, whereas the text “The notion that all Muslims are terrorists is a harmful stereotype” is an example of mention. For further contextualization of the use-mention distinction, as well as the impact of this phenomenon on downstream social bias detection, refer to Section 6. Extended analysis revealed that many of misclassifications from our detector were the result of failing to differentiate use from mention. In particular, a substantial quantity of the responses from the Llama-2 model were of the following flavor: “It is not accurate to say that {toxic_statement},” which precisely maps to a mention.
@@ 2.1 Taxonomy-Driven Synthetic Data Generation
We opted to utilize taxonomy driven generation, as this allowed us to explore a rich and structured manner through which to generate data at scale. First, we generated and iteratively updated a taxonomy of social biases. We started by prompting the Mixtral mixture-of-experts model222https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1, under greedy decoding, with the instruction: Generate a high level taxonomy for social bias that shows various groups and associated biases and stereotypes. After post-processing, we were able to arrive at a viable taxonomy with around 50 social biases across a variety of dimensions. Some examples include Anti-Asian bias, Transphobia, Ethnocentrism (the belief that one’s own culture is superior to others), etc. The full taxonomy that we utilized is available in Appendix B. Importantly, the taxonomy provides over 10 high-level topics (e.g., Racial and Ethnic Bias, Gender Bias, Classism, etc.), and a variety of sub-topics for each of these groups. In particular, we were encouraged to see that our generated taxonomy included demographic groups that are not as mainly focused on, but are still extremely relevant for the purposes hate speech detection (Nagireddy et al. 2024). One such example is Polyphobia, or prejudice against individuals who are attracted to multiple partners.
Once we generated this taxonomy, we proceeded with tailored synthetic data generation. Here, we directly leveraged the taxonomy and generated data for each leaf node (which is each sub-topic in our taxonomy). Using the prompt in Figure 1, we repeatedly instructed the aforementioned Mixtral mixture-of-experts model to generate pairs of output - jointly generating one biased and one unbiased statement. For reproducibility, we generated under nucleus sampling (Holtzman et al. 2020) with top_p at 0.95, top_k at 100, temp at 0.7, and max_new_tokens set to 1024. This method was desirable for a couple of reasons. First, the generation directly utilized the taxonomy, as each call for a generation contained both the root node (i.e., the over-arching topic) as well as the leaf node (i.e., the specific bias) directly in the context for the prompt. Next, requiring JSON formatting in the output allowed for easy post-processing of the LLM-generated text. Finally, and perhaps most interestingly, asking for paired output resulted in the generation of contrastive data. Specifically, we observed that each pair of {biased, unbiased} text precisely mapped to {use, mention} examples! Refer to Table 1 for the flavor of synthetic data obtained from this procedure.
@@ 2.2 Our Suite of Social-Bias-Detectors
social-bias refers to the models trained on the four human curated datasets mentioned in Section 2 and elaborated upon in Appendix A.1. In particular, we have three such models - each trained on top of a different base model. social-bias is trained on top of BERT333https://huggingface.co/google-bert/bert-base-uncased, social-bias-distil is trained on top of a transformer architecture (Trivedi et al. 2023) which provides most of the accuracy of a BERT-like model (Devlin et al. 2019), while being seven times faster on a CPU and two times faster on a GPU. Finally, social-bias-toxigen is trained on top of toxigen_hatebert444https://huggingface.co/tomh/toxigen˙hatebert(Hartvigsen et al. 2022). Note that these three models were trained with only human curated datasets.
@@
We hypothesized that the cascade would be a “best-of-both-worlds” approach where would do a decent job on aggregate, due to the high quality human curated datasets that it has seen. Then, to precisely combat the use-mention distinction issue, where would incorrectly flag text that is a mention as harmful, we will run any harm-labeled instances from through , a model that is specifically trained, by way of our synthetic data, to distinguish use from mention.
@@ 3.1 Evaluation Datasets and Baselines
Due to the fact that our primary axis for evaluation is the use-mention distinction, we sought to construct an evaluation set which reflected this goal. Even though datasets for such a task are not widely available (Gligoric et al. 2024), we leveraged the rich literature on counter-speech and counter-narratives in the context of hate speech. In particular, similar to (Gligoric et al. 2024), we take the Knowledge-grounded hate countering (Chung, Tekiroğlu, and Guerini 2021) and the Multi-Target CONAN (Fanton, Margherita and Bonaldi, Helena and Tekiroğlu, Serra Sinem and Guerini, Marco 2021) datasets. The knowledge-grounded dataset contains 195 hate speech and counter-narrative pairs covering multiple hate targets (islamophobia, misogyny, antisemitism, racism, and homophobia). The counter narratives are written by an expert who is tasked with composing a suitable counter-narrative response to a given hate speech using the corresponding knowledge as much as possible (Chung, Tekiroğlu, and Guerini 2021). The Multi-Target CONAN dataset consists of 5003 hate Speech and counter-narrative pairs covering multiple hate targets, including disabled, Jews, LGBT+, migrants, Muslims, people of color (POC), women. The dataset is constructed using a novel human-in-the-loop data collection methodology (Fanton, Margherita and Bonaldi, Helena and Tekiroğlu, Serra Sinem and Guerini, Marco 2021).
We combined both datasets and labeled each pair of {hate speech, counter narrative} as {bias, not bias}. In addition, we noted that a label of bias most precisely meant a label of use and similarly for not bias and mention. This is because counter-narratives are written such that they directly counteract the harmful hate speech (Gligoric et al. 2024), which results in them being excellent examples of mentions.
@@
We provide two sets of evaluations below. First, in order to compare with the results from (Gligoric et al. 2024), we utilize the same evaluation set of 180 total examples (also taken from the above two datasets), including both hate speech and counter-narratives. Second, we combine the entirety of the Knowledge-grounded hate countering and Multi-Target CONAN datasets to arrive at an evaluation set of size 555This comes from combining 195 pairs with 5003 pairs from the Knowledge-grounded hate countering and Multi-Target CONAN datasets, respectively..
@@
In addition to the reported numbers with three of the GPT models from (Gligoric et al. 2024), we also report numbers using Llama-Guard666https://huggingface.co/meta-llama/LlamaGuard-7b and Llama-Guard-2777https://huggingface.co/meta-llama/Meta-Llama-Guard-2-8B. We use these two models as both a competitive baseline as well as a point of comparison due to the size and latency of these models. To reiterate, the detectors are on the order of 100M parameters, Llama-Guard models are either 7B or 8B parameters, and GPT-4 is rumored to be orders of magnitude larger (OpenAI et al. 2024).
@@ 4 Experimental Results
In this section, we provide results on the aforementioned datasets, with all of our detectors and baselines. We report three metrics of interest: false positive rate (FPR), false negative rate (FNR), and average error rate (Avg Err) - similar to (Gligoric et al. 2024). For clarification, the average error rate is the average of the FPR and FNR. To further contextualize these metrics, a false positive refers to incorrectly providing a label of bias to benign text. Moreover, because all of the examples of benign text are examples of mentions and all the examples of harmful text are uses, a false positive represents incorrectly flagging a mention as use. Symmetrically, a false negative refers to incorrectly flagging a use as mention. For our purposes, we note that both false positives and false negatives are important. False positives represent improper moderation by flagging benign text whereas false negatives represent missed detection by failing to flag harmful text.
@@ 5.1 Results for 180 Examples Test Set
First, we comment on results for the evaluation set of 180 examples from (Gligoric et al. 2024). Despite the small size of this test, we wanted to demonstrate performance because (Gligoric et al. 2024) provided us with points of comparison to the GPT-family of models. In particular, we observe that our cascade models provide competitive performance with the GPT models. We make particular note of cascade-onetrial. Recall that the model which determines use or mention for cascade-onetrial has only seen one iteration of synthetic data generation, making it extremely desirable, as the burden of training data is substantially reduced. In absolute terms, for cascade-onetrial saw around 1K examples in training, whereas for cascade-distil saw around 190K examples in training (refer to Appendix C) for full details). Therefore, we are able to see that for this (albeit limited) set of data, the cascade approach is able to perform on par with some of the most widely used and largest models, while having a fraction of the computational cost. We do notice that both Llama-Guard models surpass even GPT-4 by a significant margin. Nevertheless, for this set of 180 examples, the cascade approach takes under 1 minute on a single A100 GPU, and is even able to be run on a CPU (taking a few hours). In comparison, access to GPT models is limited to querying via expensive external APIs, and inference with Llama-Guard models is not possible on a CPU.
@@ 5.3 Extended Analysis on Llama-Guard-2
Ignoring the increased computational overhead, we further explore the differences in performance (with respect to false positives) when comparing our detectors with the Llama-Guard models. At an instance level, we explored the differences in predictive quality between Llama-Guard and our detectors. For example, the sample Did you know that children with Down syndrome, if surrounded by regularly-developed peers, receive a fundamental boost in their development? This ”mix” fosters a positive learning atmosphere and increases engagement and attention in all students. This mean that children with Down syndrome would be a burden to society if isolated, not the contrary. was incorrectly flagged as unsafe by Llama-Guard-2, but properly classified as not_bias by our cascade-orig method. Another such sample is Even if all British Muslims had cancer, you are safe: you cannot catch it from anyone! Cancer cannot be spread, since it is a non contagious disease. Also, it is worth considering that illnesses are not connected to religions.
@@ 6 Related Work
Social bias can be defined as discrimination for, or against, a person or group, or a set of ideas or beliefs, in a way that is prejudicial or unfair (Webster et al. 2022; Bommasani and Liang 2022). Numerous studies have demonstrated that generative models exhibit undesirable behavior that amplifies social bias (Blodgett et al. 2020; Parrish et al. 2022; Smith et al. 2022; Selvam et al. 2023; Dhamala et al. 2021; Nagireddy et al. 2024). Additionally, the deficiencies of current datasets (Blodgett et al. 2021) and opacity of defining what constitutes social bias in language models and their measures (Blodgett et al. 2020; Selvam et al. 2023; Achintalwar et al. 2024) have emerged.
@@
As discussed briefly at the end of Section 1, the use-mention distinction can be thought of as the difference between using text for ill-intent and simply mentioning text without this malicious undertone. Although a subtle difference, this phenomenon is essential for many downstream applications. For example, a significant amount of content online falls under the umbrella of mentions, such as counter-speech, media reporting, education, and legal settings (Gligoric et al. 2024; Mun et al. 2023; Kirk et al. 2022; Henderson et al. 2022; Wexler, Robbennolt, and Murphy 2019). Of particular use to us is the notion of counter-speech, which refers to speech produced by users of online platforms to counteract harmful speech of others (Gligoric et al. 2024). By definition, counter-speech is an example of mention, and thus detection systems which are unable to distinguish use from mention risk contributing to downstream harm - such as improper removal of counter-speech. This in turn reduces opportunities to rectify false narratives and risks further censoring those already most affected by harmful language (Gligoric et al. 2024). Recently, there has also been work which documents the real experiences of counter-speakers and the ways in which to address existing barriers (Mun et al. 2024). An exciting by-product of our work was our own introduction to this rich space, and we believe that the area of counter-speech will continue to provide essential help in the plight of detecting and mitigating harmful content produced by generative models.
@@
There is a growing set of rich literature on guardrail models and how they can be more efficient, modular, and scalable (Achintalwar et al. 2024; Rebedea et al. 2023; Inan et al. 2023). NeMo Guardrails (Rebedea et al. 2023) is an open-source toolkit for adding modular and programmable guardrails to LLM-based conversational systems. Llama-Guard (Inan et al. 2023) is an LLM-based input-output safeguard model which has a customizable taxonomy of harms. Finally, the authors in (Achintalwar et al. 2024) go over insights from developing a host of guardrail-like models, which they refer to as detectors.
@@
Central to our work is the leveraging of capable large language models to generate synthetic data which can be used to train our guardrail models. In particular, we leverage taxonomy-guided data generation. Previous work such as GLAN (Li et al. 2024) utilizes a pre-curated taxonomy of human knowledge and capabilities as input and generates large-scale synthetic instruction data. However, as pointed out by (Sudalairaj et al. 2024), this method relies using the proprietary GPT-4 (OpenAI et al. 2024) as the teacher model, which imposes restrictions on the downstream commercial usability of the generated data. Concretely, the terms of use of proprietary models obfuscate the viability of training commercially viable models using data that is generated from proprietary or otherwise closed models - which many times forbid using their model to improve other models 999See point v under License Rights in https://llama.meta.com/llama3/license/. Thus, similar to (Sudalairaj et al. 2024), we utilize the open source Mixtral101010https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 model to generate our data. Additionally, the taxonomy guides the generation of the data, as it enables targeted coverage around the individual leaf nodes of the taxonomy (Sudalairaj et al. 2024).